29. Project Analysis and Inflation Earp Brothers, Inc., is considering investing in a machine to produce computer keyboards.*Please show all excel formulas used* Thank you!
Answer #1

Net present value is the most commonly used technique for taking capital budgeting decisions in order to determine whether the long term project is worthy of funding or not. The criteria for choosing a project by taking net present value in consideration is that the project should generate a positive net present value.

Calculation of the annual depreciation of the machine using the straight line depreciation method:-

Annual depreciation= Initial cost / Useful life

= 990,000 / 5= $198,000

Using Excel, proforma Income Statement is made:-

Book3 - Excel nitin arora x File Home Insert Page Layout Formulas Data Review View Help Tell me what you want to do A Share A

The formula sheet of the above Cash flow worksheet is given below:-

Book3 - Excel nitin arora - x File Home Insert Page Layout Formulas Data Review View Help Tell me what you want to do A Share

We can now calculate the net present value of the machine using the following formula:-

o - I Net Present value = cf, (170) + Cf₂ (1+r)? + (fn (147) where CF 011 Cash flows discount rate Initial Investment fes Sc

Calculation of Net Present Value using Excel is done below :-

Book3 - Excel nitin arora Х File Home Insert Page Layout Formulas Data Review View Help Tell me what you want to do A Share A

The formula sheet of the above NAV worksheet is given below:-

Book3 - Excel nitin arora - Х File Home Insert Page Layout Formulas Data Review View Help Tell me what you want to do A Share

The net present value of the investment is $334959.5197